# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

### v3

@@ 4.3 Math-500 Reasoning Benchmark
Reasoning is an important long-context task for LLMs. Unlike other long-context use cases, reasoning typically involves a relatively short prompt followed by a long generation, which presents unique challenges for token eviction methods. To evaluate the effectiveness of token eviction methods, we apply KeyDiff and SnapKV to the DeepSeek-R1-Distill-Qwen-7B and Llama-8B distilled models [12], and assess their performance on the Math-500 reasoning benchmark [13]. Surprisingly, we found that Llama equipped with KeyDiff and a moderate KV cache budget performs comparably to, or slightly better than, the eviction-free baseline, while also outperforming SnapKV. We kindly refer the reader to Appendix E for additional details on the reasoning task evaluation.
@@ 4.4 Ablation Study
We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff. We provide a full description of the test setup in Appendix G and summarize the findings here:
  • KeyDiff anchor choice does not greatly impact benchmark accuracies (See Table 14).
  • KeyDiff using cosine similarity as the distance metric outperforms other metrics (See Table 15)
@@ 4.5 Latency and Complexity
Additionally, in order to demonstrate that KeyDiff decreases end-to-end inference latency, we measured time to first token for the Llama 3.2 3B instruct model using different block prompt processing sizes and cache strategies. These results are visualized in Figures 17 and 7. Since KeyDiff does not require attention weight materialization, FlashAttention [8] can be used, resulting in up to 30% lower latency than TOVA and SnapKV. We compare the complexity of KeyDiff with competitors in Appendix B and perform a complete FLOP count in Section B.1.
@@ 5 Related Work
Unlike the sparse attention and KV cache compression methods, eviction methods evict KVs from the cache to reduce the size of the KV cache. As discussed in Section 2.3, the majority of the token eviction methods employ their own rules to decide the importance of the tokens by manipulating the attention score $A$. For example, by appropriately choosing the aggregation functions $\phi(A)$ of Equation 4, we can obtain existing attention-based eviction methods as discussed in Section A.1. Attention-based eviction may be a better choice when the entire prompt is being processed at once, as the eviction can be done by assessing the importance of all tokens simultaneously. However, computing the full attention score of long prompts could be prohibitively expensive in resource-constrained environments.
@@ A.1 Attention-based eviction methods
In this section, we provide a unified framework to understand prominent attention-based eviction methods. As mentioned in Equation 5, we can specify attention-based eviction methods under the unified framework with proper selection of the aggregation function $\phi(A)$ as follows:
@@ Appendix B Runtime and Memory Complexity
We analyze the runtime and memory complexity for the prominent KV cache eviction algorithms in Table 2. For a given block size $B$ and cache budget $N$, KeyDiff requires $\mathcal{O}(N+B)$ runtime and memory. The same holds true for TOVA, since it only requires computing the bottom row of $A$. Sink attention retains the $k$ first tokens in the input sequence, followed by a sliding window of size $L$, resulting in $\mathcal{O}(k+L)=O(N)$ memory and runtime, since $k+L$ equals the chosen cache budget. SnapKV computes attention over a sliding window of size $L$ against $N+B$ keys from the incoming block and the key cache, so the memory and runtime complexity is $\mathcal{O}((N+B)L)$. H2O accumulates attention weights over all tokens, and computes the attention over the current block, so it will require $\mathcal{O}(NB+B^{2})$ memory overhead and runtime. We summarize these details in Table 2
@@ B.1 FLOP count of KeyDiff
  • $\|k_{i}\|=\sqrt{k_{i}\cdot k_{i}}$: since $k_{i}\in\mathbb{R}^{d}$: $d$ multiplications, $d-1$ additions, one square root. Repeating for each $i$, this contributes $nd$ multiplications, $n(d-1)$ additions, and $n$ square roots.
@@
  • $\frac{1}{n}\sum_{i=1}^{n}c_{i}$, for $c\in\mathbb{R}^{d}$: one division, $(n-1)d$ additions.
Combining the above, we can compute the anchor vector using $2nd$ multiplications, $2nd-n-d$ additions, $n$ square roots and $n+1$ divisions.
To compute the cosine similarity score, we have from above that $\mu(\hat{K})\cdot k_{i}$ requires $nd$ multiplications and $n(d-1)$ additions. Also from above, we have that computing $\|\mu(\hat{K})\|$ requires $d$ multiplications, $d-1$ additions and one square root. We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison equals addition in cost). We can then divide through to compute $\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)}$ with $n$ divisions. Therefore, computing the cosine similarity between the anchor and each key requires $nd+d+n$ multiplications, $nd+d-1$ additions, $n+1$ divisions and one square root.
@@ C.1 Additional PCA Visualizations
In order to demonstrate the phenomena in Figure 11 persists across all heads and layers, we have included several more visualizations as seen in Figure 8, Figure 9, and Figure 10.
@@
Figure 9: Keys taken from layer $20$ and head 00 of Llama3.2-3B-Instruct
@@
Figure 10: Keys taken from layer $8$ and head $1$ of Llama3.2-3B-Instruct
@@ C.2 Derivation of KeyDiff from an Optimization Perspective
This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference. However, we can relax Section C.2 to produce a more tractable approximation to the original problem. First, we rewrite Section C.2 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
@@
To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt. We visualize the distribution of retained and evicted keys by applying sink attention [35], TOVA [24], and KeyDiff as eviction policies in Figure 11. Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed. This observation generalizes across heads and layers, as shown in Section C.1.
We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 12. This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$. The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 11. Details of how these plots were generated are discussed in Section F.1.
@@ C.3 Attention Sinks and Approximate Collinearity
At the same time, the results in [11] suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$. In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space. Our own results, seen in Figure 13 demonstrate that this is the case for both keys and queries. Our experiments show that most keys and queries lie within a small angular distance from the mean key and query. More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 14. Moreover, as seen in Figure 15, we find that the norms of keys and queries are tightly clustered around a relatively fixed value. This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction. These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query. This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries. This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in [5], however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
@@
To verify the above hypothesis, we show Figure 16 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights. This demonstrates how LLMs exploit the geometry of the hidden states and projections to limit over mixing, and selectively identify important tokens.
@@
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose $||q||=1$, the scaling parameter is $1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
@@
Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
@@
Taking the limit as $n\to\infty$ produces Lemma 3.1
@@
Note that $\alpha_{i}=\frac{y^{\top}x_{i}}{||y||||x_{i}||}=\frac{y^{\top}x_{i}}{||y||}$. If we expand $\frac{y}{||y||}$ in the basis $\{x_{1},...,x_{n}\}$ we see that
$\frac{y}{||y||}$ $=\sum\limits_{i=1}^{n}\left({\frac{y}{||y||}}^{\top}x_{i}\right)x_{i}$
@@
But then, we know that since $\left|\left|\frac{y}{||y||}\right|\right|_{2}^{2}=1$, then $\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle=1$. But we have
@@
$\bar{k}=||\bar{k}||\left(\alpha_{q}q+\sum\limits_{i=1}^{n-1}\alpha_{i}r_{i}\right)$
where $\alpha_{i}=\textrm{CosSim}(\bar{k},r_{i})$. Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$. Now we have that
$\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$ $=\frac{{k^{*}}^{\top}\left(||\bar{k}||\alpha_{q}q+\sum\limits_{i=1}^{n-1}||\bar{k}||\alpha_{i}r_{i}\right)}{||{k^{*}}||||\bar{k}||}$
$=\alpha_{q}\beta_{q}+\frac{1}{||{k^{*}}||}\sum\limits_{i=1}^{n-1}\alpha_{i}{k^{*}}^{\top}r_{i}$
@@
Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $-\frac{1}{2}$ as $\alpha\rightarrow-1$. Hence, the smaller $\textrm{CosSim}(\bar{k},q)=\alpha_{q}$ is, the smaller $\textrm{CosSim}(\bar{k},{k^{*}})$ must be.
@@ Appendix D TTFT Analysis
We have measured end-to-end inference latency (measured as time to first token (TTFT)) for Llama 3.2-3B via the standard Huggingface API when the model is using eager attention and FlashAttention as in Figure 17 and Figure 7. Tests are performed on NVIDIA A100 80GB GPUs. We test various block sizes and KV cache budgets. KeyDiff outperforms TOVA and SnapKV with FlashAttention as well as with eager attention for large cache budgets. We can see inference latency with KeyDiff is independent of block size when using FlashAttention because of KeyDiff’s linear complexity and its lack of required materialized attention weights.

### v4

@@ 4.3 Math-500 Reasoning Benchmark
Reasoning is an important long-context task for LLMs. Unlike other long-context use cases, reasoning typically involves a relatively short prompt followed by a long generation, which presents unique challenges for token eviction methods. To evaluate the effectiveness of token eviction methods, we apply KeyDiff and SnapKV to the DeepSeek-R1-Distill-Qwen-7B and Llama-8B distilled models [12], and assess their performance on the Math-500 reasoning benchmark [13]. Surprisingly, we found that Llama equipped with KeyDiff and a moderate KV cache budget performs comparably to, or slightly better than, the eviction-free baseline, while also outperforming SnapKV. We kindly refer the reader to Appendix˜E for additional details on the reasoning task evaluation.
@@ 4.4 Ablation Study
We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff. We provide a full description of the test setup in Appendix˜G and summarize the findings here:
  • KeyDiff anchor choice does not greatly impact benchmark accuracies (See Table˜15).
  • KeyDiff using cosine similarity as the distance metric outperforms other metrics (See Table˜16)
@@ 4.5 Latency and Complexity
Additionally, in order to demonstrate that KeyDiff decreases end-to-end inference latency, we measured time to first token for the Llama 3.2 3B instruct model using different block prompt processing sizes and cache strategies. These results are visualized in Figures˜17 and 7. Since KeyDiff does not require attention weight materialization, FlashAttention [8] can be used, resulting in up to 30% lower latency than TOVA and SnapKV. We compare the complexity of KeyDiff with competitors in Appendix˜B and perform a complete FLOP count in Section˜B.1.
@@ 5 Related Work
Unlike the sparse attention and KV cache compression methods, eviction methods evict KVs from the cache to reduce the size of the KV cache. As discussed in Section˜2.3, the majority of the token eviction methods employ their own rules to decide the importance of the tokens by manipulating the attention score $A$. For example, by appropriately choosing the aggregation functions $\phi(A)$ of Equation˜4, we can obtain existing attention-based eviction methods as discussed in Section˜A.1. Attention-based eviction may be a better choice when the entire prompt is being processed at once, as the eviction can be done by assessing the importance of all tokens simultaneously. However, computing the full attention score of long prompts could be prohibitively expensive in resource-constrained environments.
@@ A.1 Attention-based eviction methods
In this section, we provide a unified framework to understand prominent attention-based eviction methods. As mentioned in Equation˜5, we can specify attention-based eviction methods under the unified framework with proper selection of the aggregation function $\phi(A)$ as follows:
@@ Appendix B Runtime and Memory Complexity
We analyze the runtime and memory complexity for the prominent KV cache eviction algorithms in Table˜2. For a given block size $B$ and cache budget $N$, KeyDiff requires $\mathcal{O}(N+B)$ runtime and memory. The same holds true for TOVA, since it only requires computing the bottom row of $A$. Sink attention retains the $k$ first tokens in the input sequence, followed by a sliding window of size $L$, resulting in $\mathcal{O}(k+L)=O(N)$ memory and runtime, since $k+L$ equals the chosen cache budget. SnapKV computes attention over a sliding window of size $L$ against $N+B$ keys from the incoming block and the key cache, so the memory and runtime complexity is $\mathcal{O}((N+B)L)$. H2O accumulates attention weights over all tokens, and computes the attention over the current block, so it will require $\mathcal{O}(NB+B^{2})$ memory overhead and runtime. We summarize these details in Table˜2
@@ B.1 FLOP count of KeyDiff
  • $\|k_{i}\|=\sqrt{k_{i}\cdot k_{i}}$: since $k_{i}\in\mathbb{R}^{d}$: $d$ multiplications, d1d-1 additions, one square root. Repeating for each $i$, this contributes $nd$ multiplications, n(d1)n(d-1) additions, and $n$ square roots.
@@
  • $\frac{1}{n}\sum_{i=1}^{n}c_{i}$, for $c\in\mathbb{R}^{d}$: one division, (n1)d(n-1)d additions.
Combining the above, we can compute the anchor vector using $2nd$ multiplications, 2ndnd2nd-n-d additions, $n$ square roots and $n+1$ divisions.
[…]
@@ C.1 Additional PCA Visualizations
In order to demonstrate the phenomena in Figure˜11 persists across all heads and layers, we have included several more visualizations as seen in Figure˜8, Figure˜9, and Figure˜10.
@@
Figure 9: Keys taken from layer 2020 and head 0 of Llama3.2-3B-Instruct
@@
Figure 10: Keys taken from layer 88 and head $1$ of Llama3.2-3B-Instruct
@@ C.2 Derivation of KeyDiff from an Optimization Perspective
To leverage the observation in Section˜3.1, we minimize the sum of pairwise cosine similarities of each key retained in the cache. This can be formulated as a constrained optimization problem with the keys $K\in\mathbb{R}^{n\times d}$ whose element is $k_{i}$, and budget $N$ smaller than $n$:
@@
[…]
$|S|$ $=N$
This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference. However, we can relax Equation˜A.2 to produce a more tractable approximation to the original problem. First, we rewrite Equation˜A.2 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
@@
where $\mu(\hat{K}_{S})=\frac{1}{N}\sum\limits_{j\in S}\hat{k}_{j}$ is the empirical mean of normalized keys in $S$. This objective requires recomputing $\mu(\hat{K}_{S})$ for each candidate subset $S$. The sub-sampled mean tends to converge to the mean over the entire set, the problem can be further relaxed by replacing $\mu(\hat{K}_{S})$ with $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$. Dropping $N$ from the objective (since it doesn’t affect the solution) yields:
[…]
$|S|$ $=N$
@@
To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt. We visualize the distribution of retained and evicted keys by applying sink attention [35], TOVA [24], and KeyDiff as eviction policies in Figure˜11. Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed. This observation generalizes across heads and layers, as shown in Section˜C.1.
We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure˜12. This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$. The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure˜11. Details of how these plots were generated are discussed in Section˜F.1.
@@ C.3 Attention Sinks and Approximate Collinearity
At the same time, the results in [11] suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$. In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space. Our own results, seen in Figure˜13 demonstrate that this is the case for both keys and queries. Our experiments show that most keys and queries lie within a small angular distance from the mean key and query. More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure˜14. Moreover, as seen in Figure˜15, we find that the norms of keys and queries are tightly clustered around a relatively fixed value. This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction. These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query. This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries. This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in [5], however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
@@
To verify the above hypothesis, we show Figure˜16 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights. This demonstrates how LLMs exploit the geometry of the hidden states and projections to limit over mixing, and selectively identify important tokens.
@@
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose q=1||q||=1, the scaling parameter is $1$ and assume $k^{*}$} is a key not in $\{k_{i}\}_{i=1}^{n}$ with ${k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
@@
Note that $M\leq k_{i}^{\top}q\leq M$ and hence
@@
Taking the limit as $n\to\infty$ produces Lemma˜3.1
@@
[…]
@@
[…]
@@
[…]
where $\alpha_{i}=\textrm{CosSim}(\bar{k},r_{i})$. Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma˜C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$. Now we have that
[…]
@@
Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $\frac{1}{2}$ as $\alpha\rightarrow-1$. Hence, the smaller $\textrm{CosSim}(\bar{k},q)=\alpha_{q}$ is, the smaller $\textrm{CosSim}(\bar{k},{k^{*}})$ must be.
C.4 Correlation Analysis
We report the Spearman rank correlation between key cosine similarity and attention scores for each layer of the Llama-3.2-3B-Instruct model. Correlations are averaged over randomly sampled LongBench-Musique prompts. From Table˜3, we observe a consistently high correlation ($\rho\approx 0.94$ on average) across all layers, indicating that geometrically distinctive keys (i.e., those with low pairwise cosine similarity) are strongly aligned with tokens receiving higher attention scores. This empirical evidence supports our theoretical claim in Section˜3.3 that key diversity serves as a reliable proxy for token importance.
Table 3: Spearman correlation ($\rho$) between negative key cosine similarity and attention scores for each transformer layer of Llama-3.2-3B-Instruct.
Layer 1 2 3 4 5 6 7
$\rho$ 0.8997 0.9561 0.9332 0.9496 0.9517 0.9484 0.9587
Layer 8 9 10 11 12 13 14
$\rho$ 0.9578 0.9534 0.9570 0.9628 0.9554 0.9658 0.9578
Layer 15 16 17 18 19 20 21
$\rho$ 0.9477 0.9538 0.9280 0.9373 0.9328 0.9340 0.9311
Layer 22 23 24 Mean $\pm$ Std
$\rho$ 0.9140 0.9060 0.8950 $0.94\pm 0.02$
@@ Appendix D TTFT Analysis
We have measured end-to-end inference latency (measured as time to first token (TTFT)) for Llama 3.2-3B via the standard Huggingface API when the model is using eager attention and FlashAttention as in Figure˜17 and Figure˜7. Tests are performed on NVIDIA A100 80GB GPUs. We test various block sizes and KV cache budgets. KeyDiff outperforms TOVA and SnapKV with FlashAttention as well as with eager attention for large cache budgets. We can see inference latency with KeyDiff is independent of block size when using FlashAttention because of KeyDiff’s linear complexity and its lack of required materialized attention weights.